A rectangular garden, measuring 20m by 15 m, has a uniform strip removed from the edge of one length and the edge of one width to make a concrete walkway. If the area of the remaining garden is 204m squared , what is the width of the walkway
Answer by ptaylor(2198)


Area=Length times Width or A=LW
Let x=width of walkway
Original area is 20 times 15=300 sq meters
Area after walkway is installed is (15-x) times (20-x)=204 sq meters
(15-x)(20-x)=204
300-35x+x^2=204
x^2-35x+96=0 Quadratic in standard form: Solve using quadratic formula
x=(35+-sqrt((-35)^2-4*1*96))/2
x=(35+-sqrt(1225-384))/2
x=(35+-sqrt(841))/2
x=(35+-29)/2
x=(64)/2= 32 meters unrealistic
and
x=6/2= 3 meters answer
CK
(15-3)(20-3)=204
(12)(17)=204
204=204
Hope this helps---ptaylor